OpenStudy (anonymous):

1.4 x 10 to the 7th power/7.0 x 10 to the -2 power in scientific notation

OpenStudy (mertsj):

\[\frac{1.4 \times 10^7}{7.0 \times 10^{-2}}= \frac{1.4}{7.0} \times \frac{10^7}{10^{-2}}=.2 \times 10^9=2\times 10^8\]

OpenStudy (anonymous):

Why is it to the 9th power and not the 10th?

OpenStudy (mertsj):

Because 7 -(-2)=9

OpenStudy (anonymous):

Ok. I see. Thank you

OpenStudy (mertsj):

yw